OpenStudy (moazzam07):

Question 4: Two equations are shown: Equation 1: 3 4 (x−12)=12 Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true.

OpenStudy (anonymous):

yea?

OpenStudy (moazzam07):

3/4y−12=12

OpenStudy (moazzam07):

3/4(x-12)=12

OpenStudy (moazzam07):

i dont know how to solve it

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }y-12=12\] and\[\frac{ 3 }{ 4 }(x-12)=12\]

OpenStudy (moazzam07):

yes that

OpenStudy (moazzam07):

how do i solve it

OpenStudy (anonymous):

we add 12 to both sides for the first equation

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 y}-12+12 = 12+12\] simplifed to \[\frac{ 3 }{ 4 }y =24\]

OpenStudy (moazzam07):

yes

OpenStudy (anonymous):

multiply both sides by 4 to cancel the denominator we get 3y=96

OpenStudy (moazzam07):

ok so 3y = 96 is left

OpenStudy (anonymous):

divide both sides by 3

OpenStudy (moazzam07):

1y = 31

OpenStudy (anonymous):

32

OpenStudy (moazzam07):

y=32

OpenStudy (anonymous):

yea

OpenStudy (moazzam07):

:O ok

OpenStudy (anonymous):

2nd part try doing it the same way

OpenStudy (moazzam07):

ok

OpenStudy (moazzam07):

i ended up with 3x = 6

OpenStudy (moazzam07):

then i divided both sides by 3

OpenStudy (anonymous):

lemme show ya...

OpenStudy (moazzam07):

x=2?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 } (x-12)=12\] multiply 4 on both sides we get \[3(x-12) = 48\]

OpenStudy (moazzam07):

oh i added 12 to both sides

OpenStudy (moazzam07):

now i add 12 to both sides?

OpenStudy (anonymous):

expand \[3(x-12)\]

OpenStudy (moazzam07):

3x - 36

OpenStudy (anonymous):

3x-36=48 now add 36 to both sides

OpenStudy (moazzam07):

3x=84

OpenStudy (anonymous):

yes

OpenStudy (moazzam07):

x=28

OpenStudy (anonymous):

correct

OpenStudy (moazzam07):

OMG you are awesome thank you sooooo much

OpenStudy (anonymous):

yw

OpenStudy (moazzam07):

x=28